OpenStudy (anonymous):

STOP AND HELP PLZ! Assuming x ≠ 0 and y ≠ 0, what is the quotient of 8x^(9y³+2x^(6y²+8x³))/4y³? Which equation represents the line passing through the points (1, 4) and (-2, 13)? x - 3y = 7 3x - y = 7 x + 3y = 7 3x + y = 7 PLZ JUST HELP!

OpenStudy (zepp):

\(\huge \frac{8x^{(9y^3+2x^{(6y^2+8x^3)})}}{4y^3}\)

OpenStudy (zepp):

Like this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

plz help

OpenStudy (zepp):

Power of a power, this is an exponent rule; \(\huge \frac{8x^{(9y^3+2x^{(6y^2+8x^3)})}}{4y^3}\) would now become \(\huge \frac{8x^{(9y^3+2x)*(6y^2+8x^3)}}{4y^3}\)

OpenStudy (zepp):

\(\huge \frac{8x^{(9y^3+2x)*(6y^2+8x^3)}}{4y^3}\) Disassociate 8. \(\huge \frac{8*x^{(9y^3+2x)*(6y^2+8x^3)}}{4y^3}\) Simplify \(\huge \frac{2x^{(9y^3+2x)*(6y^2+8x^3)}}{y^3}\) We can also disassociate 2/y^3 \(\huge \frac{2}{y^3} *x^{(9y^3+2x)*(6y^2+8x^3)}\) Now multiply the stuffs in the exponent thing.

OpenStudy (anonymous):

okay

OpenStudy (zepp):

I got \(\huge \frac{2}{y^3}*x^{2y^2(27y^3+6x+36y^2+8y)}\) through simplifying.

OpenStudy (zepp):

@saifoo.khan

OpenStudy (anonymous):

okay what about the other one?